through linear substitution please help find answer to:
y=-3x-18
5x-2y=-8
and through linear elimination please help find answer to:
3x-2y=1
-6x+4y=-2
Thank you,
Lee from NC

Answer by rapaljer(4551)

Substitution method:

y=-3x-18
5x-2y=-8

Second equation:
5x - 2(______) = -8
5x - 2(-3x-18)=-8
5x + 6x + 36 = -8
11x + 36 = -8
11x + 36-36 = -8 - 36
11x = -44
x= -4

Substitute back into the first equation to find y:
y= -3x - 18
y= -3(-4) - 18
y=12 - 18 = -6

Substitute both answers into the second equation for a check:
5x-2y= -8
5(-4)-2(-6)= -8
-20+12 = -8
It checks!



Elimination method:

3x-2y=1
-6x+4y=-2
In order to eliminate the y term, multiply the first equation by 2 and add the second equation.
6x-4y = 2
-6x + 4y = -2

Adding the equations together you get 0=0 which is always true. This means that they are indeed the same line, and the solution is the entire line.

R^2 at SCC